This question deals with solving a fractional equation that is reductible to a quadratic equation. The question reads: 2/y+1 + 1/y-1 = 1
Answer by mananth(16946)

2/ y+1 +1 / y-1=1
2(y-1)+(y+1)=1
2y-2+y+1 = y^2-1
y^2-3y=0
y(y-3)=0
y=0 or y-3